OpenStudy (anonymous):

Solve the equation -1/3m-7=5

OpenStudy (hockeychick23):

so first we take -1/3m-7=5 and add 7 to both sides: -1/3m-7+=5+7 --> -1/3m=12 Now we multiply both sides by -3: (-1/3m=12)-3--> m= -36

OpenStudy (anonymous):

Thank you I understand

OpenStudy (anonymous):

Let me ask you this solve the equation 78 =2(m+3) +m. will that be-84